OpenStudy (anonymous):

\[\sqrt[3]{2y+5}+2=-3\]

OpenStudy (amistre64):

3root(2y+5) +2 = -3 ; change to exponents (2y+5)^(1/3) +2 = -3 ; -2 from each side (2y+5)^(1/3) = -5 log((2y+5)^(1/3)) = -log(5) (1/3) log(2y+5) = -log(5) log(2y+5) = -3 log(5) log(2y+5) = -log(5^3) ; unLog it. 2y+5 = -(5^3) 2y+5 = -125 2y = -125 -5 2y = -130 y = -130/2 = -65 double check with a calculator of course :) Calculator says Im right....

OpenStudy (anonymous):

hey it says solve and i got - 5 and check 2(-5)+5+2=-3

OpenStudy (amistre64):

3root(2(-5)+5) +2 <> -3 3root(2(-5)+5) +2 = 0.29 or so...